# Two-Timescale Optimization Framework for Decentralized Linear-Quadratic Optimal Control

### v1

@@ 4 Solving Inner Layer Optimization Problem
Denote
$h_{0}(X_{0})=\langle X_{0},\Omega_{k}\rangle,$
$h_{i}(X_{i})=-\left\langle X_{i},T_{2}(V^{\top}\otimes V){\rm vec}(Q)+T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]T_{1}^{\top}\Omega_{k}\right\rangle,\leavevmode\nobreak\ i=1,\dots,M,$
then optimization problem (32) is equivalent to
$\min\limits_{X_{0}\dots,X_{M}}$ $\widehat{\Theta}(X_{0},\dots,X_{M})=g(X_{0},\dots,X_{M})+\sum_{i=0}^{M}h_{i}(X_{i})$
${\rm s.t.}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak$ $X_{0}\times X_{1}\dots\times X_{M}\in\mathcal{X}=\widetilde{\Gamma}_{+}^{p}\times\widetilde{\Gamma}_{+}^{n}\times\dots\times\widetilde{\Gamma}_{+}^{n}.$
Hence, ignoring the forward sGS sweep (i.e., regard algorithm 1 as BSUM algorithm with G-S (Gauss-Seidel) rule), let
$u_{i}(X_{i},X^{k})=g(X_{1}^{k},\dots,X_{i},\dots,X_{M}^{k})+\frac{1}{2}\|X_{i}-X_{i}^{k}\|_{\mathcal{S}_{i}}^{2},\leavevmode\nobreak\ i=0,\dots,M$
with $X^{k}=(X_{0}^{k},\dots,X_{M}^{k})$. It is evident that function $u_{i}$ serves as a strictly auxiliary function, leading us to derive the subsequent theorem.
Theorem 2.
Let $\{X^{k}\}$ be the sequence generated by the BSUM algorithm with G-S rule. Then it holds
$\widehat{\Theta}(X^{k})-\widehat{\Theta}^{*}\leq\frac{c_{1}}{\sigma_{1}}\frac{1}{k},\leavevmode\nobreak\ \forall k\geq 1,$
@@
Proof.
Based on the above discussion, this theorem is a direct corollary of Theorem 1 of [24]. ∎
In algorithm 1 below, the symmetric G-S rule is employed instead of the G-S rule in order to enhance the algorithm’s performance using empirical means. Review the optimality condition of (32):
@@
00 0 $\in\partial_{X_{1}}\Theta(X_{0},X_{1},\dots,X_{M})=-H_{1}+L_{1}X_{1}+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{1}),$
@@
00 0 $\in\partial_{X_{M}}\Theta(X_{0},X_{1},\dots,X_{M})=-H_{M}+L_{M}X_{M}+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{M}),$
@@
Then, the relative residual error can be defined as
@@
${\rm err}_{X_{i}}^{k}$ $=\frac{\left\|X_{i}^{k}-\Pi_{\widetilde{\Gamma}_{+}^{n}}\left(X_{i}^{k}-L_{i}X_{i}^{k}+H_{i}\right)\right\|}{1+\|X_{i}^{k}\|+\|L_{i}X_{i}^{k}-H_{i}\|},\leavevmode\nobreak\ i=1,\dots,M.$
@@
1:Given : Initial point $(X_{0}^{0},X_{1}^{0},\dots,X_{M}^{0})$, stopping criterion parameter $\epsilon$
2:Result : ${\rm vec}(W)$
3:for $k=0,1,2,...$ do
4:     Update $\bar{X}_{M}^{k+1},\dots,\bar{X}_{1}^{k+1},X_{0}^{k+1}$ by (33), (34)
5:     Update $X_{1}^{k+1},\dots,X_{M}^{k+1}$ by (35)
6:     Compute ${\rm err}_{X_{0}}^{k+1},{\rm err}_{X_{1}}^{k+1},\dots,{\rm err}_{X_{M}}^{k+1}$ by (36)
7:     Determine ${\rm err}^{k+1}$ by (37)
8:     if ${\rm err}^{k+1}<\epsilon$ then
9:         Compute ${\rm svec}(W)$ by (38)
10:         return ${\rm vec}(W)=D_{1}{\rm svec}(W)$
11:     end if
12:end for
@@ 5 Discussion of Convergence Rate and Acceleration
1:Given : Initial point $\widetilde{W}_{0}\in\Gamma^{p},v_{0}\in\Gamma^{p},\widetilde{P}_{0}\in\mathbb{R}^{mn},w_{0}\in\mathbb{R}^{mn},\lambda_{0}\in\mathbb{R}^{mn},\theta_{0}=1,\beta_{0}>0,\gamma_{0}>0$, smoothness parameter $L_{f}$
2:Result : $\widetilde{W}_{k}$, $\widetilde{P}_{k}$
3:for $k=0,1,2,...$ do
4:     Compute $\alpha_{k}$ by (40)
5:     Compute $\eta_{g,k},\tilde{y}_{k},\tilde{\eta}_{f,k},\tilde{v}_{k}$ by (14), (15), (16), (17)
6:     Determine $u_{k}=(1+\alpha_{k})^{-1}(\widetilde{W}_{k}+\alpha_{k}v_{k}),\leavevmode\nobreak\ d_{k}=\nabla f_{1}(u_{k})+\mathcal{A}^{\top}\lambda_{k}$
7:     Update $v_{k+1}\leftarrow$Inner Layer Solver
8:     Update $\widetilde{W}_{k+1}=(1+\alpha_{k})^{-1}(\widetilde{W}_{k}+\alpha_{k}v_{k+1})$
9:     Update $\bar{\lambda}_{k+1}=\lambda_{k}+\alpha_{k}/\theta_{k}(\mathcal{A}v_{k+1}+\mathcal{B}w_{k})$
10:     Update $\widetilde{P}_{k+1}$ by (24)
11:     Update $w_{k+1}=\widetilde{P}_{k+1}+\alpha_{k}^{-1}(\widetilde{P}_{k+1}-\widetilde{P}_{k})$
12:     Update $\lambda_{k+1}=\lambda_{k}+\alpha_{k}/\theta_{k}(\mathcal{A}v_{k+1}+\mathcal{B}w_{k+1})$
13:     Update $\theta_{k+1},\gamma_{k+1},\beta_{k+1}$ by (18)
14:     if Stopping Criterion==absent$=$ =True then
15:         return $\widetilde{W}_{k+1},\widetilde{P}_{k+1}$
16:     end if
17:end for
@@
$\mathcal{E}_{k}=\mathcal{L}(\widetilde{W}_{k},\widetilde{P}_{k},\lambda^{*})-\mathcal{L}(\widetilde{W}^{*},\widetilde{P}^{*},\lambda_{k})+\frac{\gamma_{k}}{2}\|v_{k}-\widetilde{W}^{*}\|^{2}+\frac{\beta_{k}}{2}\|w_{k}-\widetilde{P}^{*}\|^{2}+\frac{\theta_{k}}{2}\|\lambda_{k}-\lambda^{*}\|^{2},$
where $(\widetilde{W}^{*},\widetilde{P}^{*},\lambda^{*})$ is a saddle point of $\mathcal{L}$, i.e.,
$\mathcal{L}(\widetilde{W}^{*},\widetilde{P}^{*},\lambda)\leq\mathcal{L}(\widetilde{W}^{*},\widetilde{P}^{*},\lambda^{*})\leq\mathcal{L}(\widetilde{W},\widetilde{P},\lambda^{*}),\leavevmode\nobreak\ \forall(\widetilde{W},\widetilde{P},\lambda).$
The convergence analysis of Algorithm 2 can be derived, as stated in the following theorem.
Theorem 3.
Under the initial condition
$\gamma_{0}>0,\leavevmode\nobreak\ \leavevmode\nobreak\ \text{\rm and }\beta_{0}>0$
and the condition
@@
$\left\|\mathcal{A}\widetilde{W}_{k}+\mathcal{B}\widetilde{P}_{k}\right\|\leq\theta_{k}\mathcal{R}_{0},\quad\left|F(\widetilde{W}_{k},\widetilde{P}_{k})-F^{*}\right|\leq\theta_{k}\left(\mathcal{E}_{0}+\left\|\lambda^{*}\right\|\mathcal{R}_{0}\right),$
@@
Proof.
Noticing that $f_{1}(\widetilde{W})$ is a linear function, $\nabla f_{1}$ is Lipschitz continuous with Lipschitz parameter $L_{f_{1}}=0$. Then, this theorem is a direct corollary of Theorem 4.1 of [31]. ∎
The subsequent proposition elucidates the asymptotic behavior of optimization problem (8) when the parameter $\gamma$ approaches zero. Therefore, it is possible to achieve a near-optimal centralized controller by selecting a sufficiently small value for $\gamma$ and utilizing Algorithm 2.
@@
Assume that $C^{\top}C\succ 0$, and denote
@@
$F(\widetilde{W})$ $=f(\widetilde{W})+\delta_{\widehat{\mathcal{A}}\widetilde{W}=0}(\widetilde{W}),$
where $\widehat{\mathcal{A}}$ is defined in (46). Then it holds
$\inf F^{\gamma_{n}}\to\inf F$
Proof.
Obviously, $F^{\gamma_{n}}(\widetilde{W})\stackrel{{\scriptstyle\rm e}}{{\longrightarrow}}F(\widetilde{W})$ (the epigraph convergence given below in Section 6). Based on the fact
@@
$F(\widetilde{W})$ is level-bounded. Additionally, all $F^{\gamma_{n}}$ are closed convex function, and thus all the sets ${\rm lev}_{\leq\alpha}F^{\gamma_{n}}$ is connected. Then, $\{F^{\gamma_{n}}\}_{n\in\mathbb{N}}$ is eventually level-bounded by 7.32 (c) of [41]. Hence
$\inf F^{\gamma_{n}}\to\inf F$
by Theorem 7.33 of [41]. ∎
Above, our choice of sparsity promoting function is the weighted-$l_{1}$ norm, i.e., $g(P)=\sum_{i,j}w_{ij}|P_{ij}|$, which is a convex but not a strongly convex function. Hence, regretfully, $\mu_{g}=0$, and algorithm (13) exhibits $\mathcal{O}\left(\frac{1}{k}\right)$ convergence rate. Alternatively, designing a strongly convex sparsity promoting function is one approach to resolving the issue described above. For instance, let $g_{2}(x):\mathbb{R}\to\mathbb{R}$ be
$g_{2}(x)=\left\{\begin{aligned} &\frac{1}{2}a_{1}x^{2}+b_{1}x,\leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ x\leq 0,\\ &\frac{1}{2}a_{2}x^{2}+b_{2}x,\leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ x>0,\end{aligned}\right.$ (41)
@@
$a_{1}\geq 0,\leavevmode\nobreak\ a_{2}\geq 0,\leavevmode\nobreak\ b_{1}\leq 0\leq b_{2},\leavevmode\nobreak\ {\rm and}\leavevmode\nobreak\ b_{2}-b_{1}>0.$
Obviously, by choosing $a_{1},a_{2}>0,b_{1}<0<b_{2}$, $g_{\mathcal{Q}}(P)$ becomes a strongly convex sparsity promoting function. Hence, there comes a surprising result that designing an algorithm with an accelerated convergence rate is equivalent to designing a strongly convex sparsity promoting function. If we consider the piecewise quadratic function $g_{\mathcal{Q}}(P)$, the primal optimization problem becomes
$\min_{\widetilde{W},\widetilde{P}}\leavevmode\nobreak\ \leavevmode\nobreak$ $f(\widetilde{W})+\gamma g_{\mathcal{Q}}(\widetilde{P})$
${\rm s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $\mathcal{A}\widetilde{W}+\mathcal{B}\widetilde{P}=0.$
@@
$=\left(\underset{P_{ij}\in\mathbb{R}}{\operatorname*{argmin}}\left\{\sum_{i,j}\left(\gamma w_{ij}g_{2}(P_{ij})+\frac{\rho_{k}}{2}\left(P_{ij}-\Omega^{k}_{ij}\right)^{2}\right)\right\}\right)_{mn\times 1},$
where $\Omega^{k}$ is given by (23). Focus on the following subproblem
$\min_{P_{ij}\in\mathbb{R}}\gamma w_{ij}g_{2}(P_{ij})+\frac{\rho_{k}}{2}\left(x-\Omega^{k}_{ij}\right)^{2}$ (42)
with
$g_{2}(P_{ij})=\left\{\begin{aligned} &\frac{1}{2}a_{1}P_{ij}^{2}+b_{1}P_{ij},\leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ P_{ij}\leq 0\\ &\frac{1}{2}a_{2}P_{ij}^{2}+b_{2}P_{ij},\leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ P_{ij}>0\end{aligned}\right.$
given by (41), and let $P_{ij}^{*}$ be the global minimizer of optimization problem (42).
Theorem 5.
Provided that $a_{1}>0,a_{2}>0,b_{1}<0<b_{2}$, then the existence and uniqueness of $P_{ij}^{*}$ holds. Moreover,
$P_{ij}^{*}=\left\{\begin{aligned} &0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ \Omega_{ij}^{k}<\frac{\gamma w_{ij}b_{2}}{\rho_{k}},\Omega_{ij}^{k}\geq 0,\\ &0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ \Omega_{ij}^{k}>\frac{\gamma w_{ij}b_{1}}{\rho_{k}},\Omega_{ij}^{k}<0,\\ &\frac{\rho_{k}\Omega_{ij}^{k}-\gamma w_{ij}b_{2}}{\gamma w_{ij}a_{2}+\rho_{k}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ \Omega_{ij}^{k}\geq\frac{\gamma w_{ij}b_{2}}{\rho_{k}},\Omega_{ij}^{k}\geq 0,\\ &\frac{\rho_{k}\Omega_{ij}^{k}-\gamma w_{ij}b_{1}}{\gamma w_{ij}a_{1}+\rho_{k}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ \Omega_{ij}^{k}\leq\frac{\gamma w_{ij}b_{1}}{\rho_{k}},\Omega_{ij}^{k}<0.\end{aligned}\right.$ (43)
Proof.
@@
it is obvious that optimization problem (42) is strongly convex. Thus, $P_{ij}^{*}$ exists and is unique. Denote $\varphi(P_{ij})=\gamma w_{ij}g_{2}(P_{ij})+\frac{\rho_{k}}{2}\left(P_{ij}-\Omega^{k}_{ij}\right)^{2}$. When $\Omega_{ij}^{k}\geq 0$, we can obtain
$\varphi(P_{ij})=\left\{\begin{aligned} &\left(\frac{\gamma w_{ij}a_{1}+\rho_{k}}{2}\right)P_{ij}^{2}+\left(\gamma w_{ij}b_{1}-\rho_{k}\Omega_{ij}^{k}\right)P_{ij}+\frac{\rho_{k}}{2}{\Omega_{ij}^{k}}^{2},\leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ P_{ij}\leq 0,\\ &\left(\frac{\gamma w_{ij}a_{2}+\rho_{k}}{2}\right)P_{ij}^{2}+\left(\gamma w_{ij}b_{2}-\rho_{k}\Omega_{ij}^{k}\right)P_{ij}+\frac{\rho_{k}}{2}{\Omega_{ij}^{k}}^{2},\leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ P_{ij}>0.\end{aligned}\right.$
@@
$P_{ij}^{*}=\left\{\begin{aligned} &0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ \Omega_{ij}^{k}<\frac{\gamma w_{ij}b_{2}}{\rho_{k}},\\ &\frac{\rho_{k}\Omega_{ij}^{k}-\gamma w_{ij}b_{2}}{\gamma w_{ij}a_{2}+\rho_{k}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ \Omega_{ij}^{k}\geq\frac{\gamma w_{ij}b_{2}}{\rho_{k}}.\end{aligned}\right.$

### v3

@@ 4 Solving Inner Layer Optimization Problem
Corollary 1.
Let $\{X^{k}\}$ be the sequence generated by the BSUM algorithm with G-S rule for solving optimization problem (38). Then it holds
$\widehat{\Theta}(X^{k})-\widehat{\Theta}^{*}\leq\frac{c_{1}}{\sigma_{1}}\frac{1}{k},~{}\forall k\geq 1,$
@@
[…]
In Algorithm 2 below, the sGS rule is employed instead of the G-S rule in order to enhance the algorithm’s performance using empirical means. Review the optimality condition of (38):
@@
00 0 $\in\partial_{X_{1}}\Theta(X_{0},\dots,X_{M})=-H_{1}+L_{1}X_{1}+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{1}),$
@@
00 0 $\in\partial_{X_{M}}\Theta(X_{0},\dots,X_{M})=-H_{M}+L_{M}X_{M}+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{M}),$
@@
Then, for $i=1,\dots,M$, the relative residual error can be defined as
@@
${\rm err}_{X_{i}}^{k}$ $=\frac{\left\|X_{i}^{k}-\Pi_{\widetilde{\Gamma}_{+}^{n}}\left(X_{i}^{k}-L_{i}X_{i}^{k}+H_{i}\right)\right\|}{1+\|X_{i}^{k}\|+\|L_{i}X_{i}^{k}-H_{i}\|}.$
@@
It is noteworthy that, for $i=0,\dots,M$, the auxiliary function $u_{i}$ is nothing but the nonseparable term $g$ of dual function $\Theta$ plus the special proximal term $\|X_{i}-X_{i}^{k}\|_{\mathcal{S}_{i}}^{2}$ designed in this paper. Hence, in fact, the design of proximal terms is equivalent to the design of the auxiliary functions of BSUM framework. Significantly, we have so far designed the special sequence of auxiliary functions $\{u_{i}\}_{i=0}^{M}$ such that problem (39) admits a closed-form solution, and thus decreased the algorithmic complexity of Inner Layer Solver; see Remark 11 and Algorithm 2 below.
@@
  Given : Initial point $(X_{0}^{0},X_{1}^{0},\dots,X_{M}^{0})$, stopping criterion parameter $\epsilon$
  Result : ${\rm vec}(W)$
  for $k=0,1,2,...$ do
     Update $\bar{X}_{M}^{k+1},\dots,\bar{X}_{1}^{k+1},X_{0}^{k+1}$ by (40), (42)
     Update $X_{1}^{k+1},\dots,X_{M}^{k+1}$ by (43)
     Compute ${\rm err}_{X_{0}}^{k+1},{\rm err}_{X_{1}}^{k+1},\dots,{\rm err}_{X_{M}}^{k+1}$ by (45)
     Determine ${\rm err}^{k+1}$ by (46)
     if ${\rm err}^{k+1}<\epsilon$ then
        Compute ${\rm svec}(W)$ by (47)
        return  ${\rm vec}(W)=D_{1}{\rm svec}(W)$
     end if
  end for
@@ 5 Discussion of Convergence Rate and Acceleration
In this section, the convergence of Algorithm 3 (using iteration scheme (21) for solving problem (15)) is studied. Additionally, the optimization problem induced by quadratic sparsity promoting function is of concern, which is shown to exhibit an accelerated convergence rate.
@@
  Given : Initial point $\widetilde{W}_{0}\in\Gamma^{p},v_{0}\in\Gamma^{p},\widetilde{P}_{0}\in\mathbb{R}^{mn},w_{0}\in\mathbb{R}^{mn},\lambda_{0}\in\mathbb{R}^{mn},\theta_{0}=1,\beta_{0}>0,\gamma_{0}>0$, smoothness parameter $L_{f}$
  Result : $\widetilde{W}_{k}$, $\widetilde{P}_{k}$
  for $k=0,1,2,...$ do
     Compute $\alpha_{k}$ by (49)
     Compute $\eta_{g,k},\widetilde{y}_{k},\widetilde{\eta}_{f,k},\widetilde{v}_{k}$ by (22), (23), (24), (25)
     Determine $u_{k}=(1+\alpha_{k})^{-1}(\widetilde{W}_{k}+\alpha_{k}v_{k}),~{}d_{k}=\nabla f_{1}(u_{k})+\mathcal{A}^{\top}\lambda_{k}$
     Update $v_{k+1}\leftarrow$Inner Layer Solver
     Update $\widetilde{W}_{k+1}=(1+\alpha_{k})^{-1}(\widetilde{W}_{k}+\alpha_{k}v_{k+1})$
     Update $\bar{\lambda}_{k+1}=\lambda_{k}+\alpha_{k}/\theta_{k}(\mathcal{A}v_{k+1}+\mathcal{B}w_{k})$
     Update $\widetilde{P}_{k+1}$ by (31)
     Update $w_{k+1}=\widetilde{P}_{k+1}+\alpha_{k}^{-1}(\widetilde{P}_{k+1}-\widetilde{P}_{k})$
     Update $\lambda_{k+1}=\lambda_{k}+\alpha_{k}/\theta_{k}(\mathcal{A}v_{k+1}+\mathcal{B}w_{k+1})$
     Update $\theta_{k+1},\gamma_{k+1},\beta_{k+1}$ by (26)
     if Stopping Criterion==absent$=$ =True then
        return  $\widetilde{W}_{k+1},\widetilde{P}_{k+1}$
     end if
  end for
Remark 14.
Denoting
$r_{k+1}=\mathcal{A}\widetilde{W}_{k+1}+\mathcal{B}\widetilde{P}_{k+1},$
$s_{k+1}=\rho\mathcal{A}^{\top}\mathcal{B}(\widetilde{P}_{k+1}-\widetilde{P}_{k})$
with constant $\rho>0$, it is recommended that the stopping criterion of Algorithm 3 can be selected as
$\|r_{k}\|\leq\epsilon^{(k)}_{\rm pri},~{}\|s_{k}\|\leq\epsilon^{(k)}_{\rm dua},$
where
$\epsilon^{(k)}_{\rm pri}=\sqrt{M+mn}\epsilon_{1}+\epsilon_{2}\max\{\|\mathcal{A}\widetilde{W}_{k}\|,\|\mathcal{B}\widetilde{P}_{k}\|\},$
$\epsilon^{(k)}_{\rm dua}=p\epsilon_{1}+\epsilon_{2}\|\mathcal{A}^{\top}\lambda_{k}\|$
with some constants $\epsilon_{1},\epsilon_{2}$ [54]. Moreover, the stopping criterion can be selected in many other ways; see [55] for some examples.
@@
$\mathcal{E}_{k}=$ $\mathcal{L}(\widetilde{W}_{k},\widetilde{P}_{k},\lambda^{*})-\mathcal{L}(\widetilde{W}^{*},\widetilde{P}^{*},\lambda_{k})+\frac{\gamma_{k}}{2}\|v_{k}-\widetilde{W}^{*}\|^{2}$
$+\frac{\beta_{k}}{2}\|w_{k}-\widetilde{P}^{*}\|^{2}+\frac{\theta_{k}}{2}\|\lambda_{k}-\lambda^{*}\|^{2},$
where $(\widetilde{W}^{*},\widetilde{P}^{*},\lambda^{*})$ is a saddle point of $\mathcal{L}$.
$\mathcal{L}(\widetilde{W}^{*},\widetilde{P}^{*},\lambda)\leq\mathcal{L}(\widetilde{W}^{*},\widetilde{P}^{*},\lambda^{*})\leq\mathcal{L}(\widetilde{W},\widetilde{P},\lambda^{*}),~{}\forall(\widetilde{W},\widetilde{P},\lambda).$
The convergence analysis of Algorithm 3 can be derived, as stated in the following corollary.
Corollary 2.
Under the initial condition $\gamma_{0}>0$, $\beta_{0}>0$ and the condition
@@
$\left\|\mathcal{A}\widetilde{W}_{k}+\mathcal{B}\widetilde{P}_{k}\right\|\leq\theta_{k}\mathcal{R}_{0},$
$\left|F(\widetilde{W}_{k},\widetilde{P}_{k})-F^{*}\right|\leq\theta_{k}\left(\mathcal{E}_{0}+\left\|\lambda^{*}\right\|\mathcal{R}_{0}\right),$
@@
[…]
The following proposition elucidates the asymptotic behavior of optimization problem (15) when the parameter $\gamma_{n}\to 0$. Therefore, it is possible to achieve a near-optimal centralized controller by selecting a sufficiently small value for $\gamma$ and utilizing Algorithm 3.
@@
Let $\gamma_{n}\to 0$. Assume that $C^{\top}C\succ 0$, and denote
@@
$F(\widetilde{W})$ $=f(\widetilde{W})+\delta_{\widehat{\mathcal{A}}\widetilde{W}=0}(\widetilde{W}).$
Then it holds $\inf F^{\gamma_{n}}\to\inf F.$
Proof: Obviously, $F^{\gamma_{n}}(\widetilde{W})\stackrel{{\scriptstyle\rm e}}{{\longrightarrow}}F(\widetilde{W})$ (the epigraph convergence; see Definition 7.1 of [56]). Based on the fact
@@
[…]
Above, our choice of sparsity promoting function is the weighted-$\ell_{1}$-norm, i.e., $g(P)=\sum_{i,j}w_{ij}|P_{ij}|$, which is convex but not strongly convex. Hence, regretfully, algorithm (21) exhibits $\mathcal{O}\left(\frac{1}{k}\right)$ convergence rate. Alternatively, designing a strongly convex sparsity promoting function is one approach to solving the issue described above. For instance, let $g_{2}:\mathbb{R}\to\mathbb{R}$ be
$g_{2}(x)=\left\{\begin{aligned} &\frac{1}{2}a_{1}x^{2}+b_{1}x,~{}{\rm if}~{}x\leq 0,\\ &\frac{1}{2}a_{2}x^{2}+b_{2}x,~{}{\rm if}~{}x>0,\end{aligned}\right.$ (50)
@@
$a_{1}\geq 0,~{}a_{2}\geq 0,~{}b_{1}\leq 0\leq b_{2},~{}{\rm and}~{}b_{2}-b_{1}>0.$
Obviously, by choosing $a_{1},a_{2}>0,b_{1}<0<b_{2}$, $g_{\mathcal{Q}}(P)$ becomes a strongly convex sparsity promoting function with modulus
$\mu_{g_{\mathcal{Q}}}=\min\limits_{i,j}\left\{w_{ij}\min\{a_{1},a_{2}\}\right\}.$
Hence, there comes a surprising result that designing an algorithm with an accelerated convergence rate is equivalent to designing a strongly convex sparsity promoting function. If we consider the piecewise quadratic function $g_{\mathcal{Q}}(P)$, the primal optimization problem becomes
$\min_{\widetilde{W},\widetilde{P}}{}{}$ $f(\widetilde{W})+\gamma g_{\mathcal{Q}}(\widetilde{P})$
${\rm s.t.}{}{}$ $\mathcal{A}\widetilde{W}+\mathcal{B}\widetilde{P}=0.$
@@
$\widetilde{P}_{k+1}=\mathbf{prox}_{\tau_{k}\gamma g_{\mathcal{Q}}}^{\mathbb{R}^{mn\times 1}}(\widetilde{P}_{k}+\frac{\alpha_{k}\beta_{k}}{\eta_{g,k}}(w_{k}-\widetilde{P}_{k})-\tau_{k}\mathcal{B}^{\top}\bar{\lambda}_{k+1}).$
Remarkably, $\widetilde{P}_{k+1}$ can be obtained by the following theorem.
Theorem 3.
Provided that $a_{1}>0,a_{2}>0,b_{1}<0<b_{2}$, letting
$\hat{P}_{ij;k+1}^{*}=\left\{\begin{aligned} &0,~{}{\rm if}~{}\Omega_{ij}^{k}<\frac{\gamma w_{ij}b_{2}}{\rho_{k}},\Omega_{ij}^{k}\geq 0,\\ &0,~{}{\rm if}~{}\Omega_{ij}^{k}>\frac{\gamma w_{ij}b_{1}}{\rho_{k}},\Omega_{ij}^{k}<0,\\ &\frac{\rho_{k}\Omega_{ij}^{k}-\gamma w_{ij}b_{2}}{\gamma w_{ij}a_{2}+\rho_{k}},~{}{\rm if}~{}\Omega_{ij}^{k}\geq\frac{\gamma w_{ij}b_{2}}{\rho_{k}},\Omega_{ij}^{k}\geq 0,\\ &\frac{\rho_{k}\Omega_{ij}^{k}-\gamma w_{ij}b_{1}}{\gamma w_{ij}a_{1}+\rho_{k}},~{}{\rm if}~{}\Omega_{ij}^{k}\leq\frac{\gamma w_{ij}b_{1}}{\rho_{k}},\Omega_{ij}^{k}<0,\end{aligned}\right.$
where $\Omega^{k}$ is given in (30), we have
$\widetilde{P}_{k+1}=(\hat{P}_{11;k+1}^{*},\hat{P}_{21;k+1}^{*},\dots,\hat{P}_{mn;k+1}^{*})^{\top}.$ (51)
Proof: By the definition of $\widetilde{P}_{k+1}$, we have
@@
$=\left(\underset{P_{ij}\in\mathbb{R}}{\operatorname*{argmin}}\left\{\gamma w_{ij}g_{2}(P_{ij})+\frac{\rho_{k}}{2}\left(P_{ij}-\Omega^{k}_{ij}\right)^{2}\right\}\right)_{mn},$
@@
it is obvious that optimization problem
$\hat{P}_{ij;k+1}^{*}\in\min_{P_{ij}\in\mathbb{R}}\gamma w_{ij}g_{2}(P_{ij})+\frac{\rho_{k}}{2}\left(x-\Omega^{k}_{ij}\right)^{2}$ (52)
is strongly convex. Thus, $\hat{P}_{ij;k+1}^{*}$ exists and is unique. Denote $\varphi(P_{ij})=\gamma w_{ij}g_{2}(P_{ij})+\frac{\rho_{k}}{2}\left(P_{ij}-\Omega^{k}_{ij}\right)^{2}$. When $\Omega_{ij}^{k}\geq 0$, we can obtain
$\varphi(P_{ij})=\left\{\begin{aligned} &\left(\frac{\gamma w_{ij}a_{1}+\rho_{k}}{2}\right)P_{ij}^{2}+\left(\gamma w_{ij}b_{1}-\rho_{k}\Omega_{ij}^{k}\right)P_{ij}\\ &\qquad\qquad\qquad+\frac{\rho_{k}}{2}{\Omega_{ij}^{k}}^{2},~{}{\rm if}~{}P_{ij}\leq 0,\\ &\left(\frac{\gamma w_{ij}a_{2}+\rho_{k}}{2}\right)P_{ij}^{2}+\left(\gamma w_{ij}b_{2}-\rho_{k}\Omega_{ij}^{k}\right)P_{ij}\\ &\qquad\qquad\qquad+\frac{\rho_{k}}{2}{\Omega_{ij}^{k}}^{2},~{}{\rm if}~{}P_{ij}>0.\end{aligned}\right.$
@@
$\hat{P}_{ij;k+1}^{*}=\left\{\begin{aligned} &0,~{}~{}~{}~{}{\rm if}~{}\Omega_{ij}^{k}<\frac{\gamma w_{ij}b_{2}}{\rho_{k}},\\ &\frac{\rho_{k}\Omega_{ij}^{k}-\gamma w_{ij}b_{2}}{\gamma w_{ij}a_{2}+\rho_{k}},~{}~{}~{}~{}{\rm if}~{}\Omega_{ij}^{k}\geq\frac{\gamma w_{ij}b_{2}}{\rho_{k}}.\end{aligned}\right.$